OpenStudy (anonymous):

y=-1/2x+3 what is the slope: what is the y-intercept: i got -1/2x as slope and 3 for y- intercept is this correct?

OpenStudy (anonymous):

YES! that is right

OpenStudy (anonymous):

thank you i have one more to check as well

OpenStudy (anonymous):

y=-x+1 and y=x-5 is this perpendicular or neither? i said perpindicular is this correct?

OpenStudy (anonymous):

um, i believe they are neither. for something to be perpendicular, it has to have the negative recipical of the other number. for example, -1/3 is the recepical of 3. the recipical of your question would be -1/x

OpenStudy (anonymous):

oh okay thank you for all your help.

OpenStudy (anonymous):

sure :) happy to help!